OpenStudy (anonymous):

x^2-5x+k, k>0 find all the values for k so that the trinomial can be factored using integers

OpenStudy (anonymous):

let x^2-5x+k=(x+m)(x+n) this gives m+n=-5 and mn=k; as k>0. so both m and n must be either +ve or -ve. but +ve m,n is not possible as m+n=-5; so m,n is -ve, (m,n)=(-1,-4) or( -2,-3) or (-3,-2) or (-4,-1); this gives k={4,6}